# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v1

@@ Appendix B Closed-form expressions of the hypergeometric model
In Euclidean spaces of odd dimension. When $d$ is an odd integer, $\kappa=k$ is a non-negative integer and $\mu$ is a positive integer, the hypergeometric model yields a truncated polynomial (Emery and Alegría,, 2022, Section 2.5.3). Specifically, if $\mu$ is odd, then
@@
The cases $k=0$, $\mu=1$ and $d=1$, $3$ and $5$ correspond to the well-known triangular, spherical and pentaspherical models (Matérn,, 1986), which behave linearly near the origin:
@@
whereas the cases $d=3$, $\mu=1$ and $k=1$ or $2$ correspond to the so-called cubic and penta models (Chilès and Delfiner,, 2012), which are associated with Gaussian RFs that are once and twice mean-squared differentiable, respectively:
@@
    When $\kappa=0$, formula 7.3.2.92 of Prudnikov et al., (1990) gives:
@@
    which corresponds to the circular correlation model (Chilès and Delfiner,, 2012, 2.46), which behaves linearly near the origin.
@@
    When $\kappa=\frac{1}{2}$, formula 7.3.2.99 of Prudnikov et al., (1990) gives:
@@
    which is the montée (upgrading) (Matheron,, 1965, I.4.18) of order $1$ of the spherical correlation model in $\mathbb{R}^{3}$.
@@
    When $\kappa=1$, one obtains (Weisstein,, 2025):

### v2

@@ Appendix B Closed-form expressions of the hypergeometric model
In Euclidean spaces of odd dimension. When $d$ is an odd integer, $\kappa=k$ is a nonnegative integer and $\mu$ is a positive integer, the hypergeometric model yields a truncated polynomial (Emery and Alegría, 2022, Section 2.5.3). Specifically, if $\mu$ is odd, then
@@
The cases $k=0$, $\mu=1$ and $d=1$, 33 and $5$ correspond to the well-known triangular, spherical and pentaspherical models (Matérn, 1986), which behave linearly near the origin:
@@
whereas the cases $d=3$, $\mu=1$ and $k=1$ or $2$ correspond to the so-called cubic and penta models (Chilès and Delfiner, 2012), which are associated with Gaussian RFs that are once and twice mean-squared differentiable, respectively:
@@
    When $\kappa=0$, formula 7.3.2.92 of Prudnikov et al. (1990) gives:
@@
    which corresponds to the circular correlation model (Chilès and Delfiner, 2012, 2.46), which behaves linearly near the origin.
@@
    When $\kappa=\frac{1}{2}$, formula 7.3.2.99 of Prudnikov et al. (1990) gives:
@@
    which is the montée (upgrading) (Matheron, 1965, I.4.18) of order $1$ of the spherical correlation model in $\mathbb{R}^{3}$.
@@
    When $\kappa=1$, one obtains (Weisstein, 2025):